In Millikan's experiment, the charge on each drop of oil was measured in coulombs. Imagine the same experiment, but with charges measured in a fictitious unit called a zeet (Z).

Drop Charge (Z)
A 2.25×10−14
B 4.05×10−14
C 4.50×10−14
D 4.95×10−14
E 6.75×10−14

What is the charge on an electron in zeets?
Answer #1

We can determine the charge on the electron in zeets, by finding the greatest coom factor among the measured charges and their differences. The differences can be calculated by subtracting each of the values from the previous value,

B-A = 4.05×10−14-2.25×10−14 = 1.85×10−14

C-B =4.50×10−14 -4.05×10−14= 0.45×10−14

D-C = 4.95×10−14-4.50×10−14= 0.45×10−14

E-D = 6.75×10−14-4.95×10−14 = 1.80×10−14

The greatest common factor is 0.45×10−14, so the charge on the electron in zeets is 0.45×10−14zeets.